At 2000 °C the equilibrium constant for the reaction below is Kc= 2.4x103 . If the initial concentration of NO is 0.500 M, what are the equilibrium concentrations of each substance?

2 NO (g)    N2 (g) + O2 (g)
Answer #1

Solution:-

NO (g) N2 (g) O2 (g)
I   0.500   0 0
C   -2x   +x   +x
E   (0.500 - 2x)   x x

Now,

Kc =  [N2] [O2] / [NO]2

(2.4 × 103) = (x.x) / (0.500 - 2x)2

(2.4 × 103) = x2 / (0.500 - 2x)2

On taking square roots on both sides, we get

48.99 = x / (0.500 - 2x)

24.49 - 97.98 x = x

24.49 = 98.98 x

x = (24.49) / (98.98) = 0.2474 M

x = [N2] = [O2] = 0.2474 M

[NO] = (0.500 - 2x) = [0.500 - 2 (0.2474)] = 0.0052 M

--------------------------------------------------------------------